In a 30 60 90 triangle the shorter leg is 8.5 i need to find x and y
Answer by stanbon(75887)

In a 30 60 90 triangle the shorter leg is 8.5 i need to find x and y
---
The shorter leg is (1/2) of the hypotenuse.
So the hypotenuse is 2(8.5) = 17
---
Use Pythagoras to find the langer leg:
x^2 + 8.5^2 = 17^2
x^2 = 216.75
x = 14.72
=========================
Cheers,
Stan H.